# Robust Max Statistics for High-Dimensional Inference

### v1

@@ Appendix F Results on median-of-means estimators
Proof. The proof of Lemma 12 can be repeated with the i.i.d. random variables $\frac{1}{2}(X_{ij}-X_{i^{\prime}j})^{2}$, playing the role that $X_{ij}$ previously did. Also note that because $\textup{var}\big{(}\frac{1}{2}(X_{ij}-X_{i^{\prime}j})^{2}\big{)}\lesssim\sigma_{j}^{4}$ holds under Assumption 1(i), the parameter $\sigma_{j}^{4}$ plays the role that $\sigma_{j}^{2}$ did in the context of Lemma 12. ∎
@@
Since the independent random variables $X_{1j}/\sigma_{j}$ and $X_{(1+\ell_{n}/2)j}/\sigma_{j}$ have densities whose $L^{\infty}$ norms are $\mathcal{O}(1)$ under Assumption 1(ii), it follows from Young’s convolution inequality [53, p.178] that the random variable $\frac{1}{\sqrt{2}\sigma_{j}}(X_{1j}-X_{(\ell_{n}/2+1)j})$ also has a density whose $L^{\infty}$ norm is $\mathcal{O}(1)$, and so
@@
where the last step uses $\ell_{n}\asymp n/\log(n)$. Note also that the the final bound $n^{-(2+\delta)}$ can be replaced with any fixed positive power of $n^{-1}$, but the current form is all that is needed. ∎
@@ Appendix G Proof of Proposition 1
Elliptical case. Suppose $X_{1}$ is a centered elliptical random vector of the form $X_{1}=\eta_{1}\Sigma^{1/2}Z_{1}/\|Z_{1}\|_{2}$, where $Z_{1}$ is a standard Gaussian $p$-dimensional Gaussian vector, and $\eta_{1}$ is independent of $Z_{1}$ with $\mathbf{E}(\eta_{1}^{2})=p$. We first check the $L^{q}$-$L^{2}$ moment equivalence condition (i). Letting $w=\Sigma^{1/2}v$ for a generic vector $v\in\mathbb{R}^{p}$, a direct calculation gives
@@
Because the distribution of $U_{1}=Z_{1}/\|Z_{1}\|_{2}$ is invariant to orthogonal transformations, it follows that the random variables $\langle w,U_{1}\rangle$ and $\|w\|_{2}\langle e_{1},U_{1}\rangle$ are equal in distribution, where $e_{1}=(1,0,\dots,0)$. Therefore,
@@
The quantity $\|U_{11}\|_{L^{q}}$ is at most of order $1/\sqrt{p}$, which can be shown as follows. Due to the the independence of $U_{11}$ and $\|Z_{1}\|_{2}$, we have $\|Z_{11}\|_{L^{q}}=\|\|Z_{1}\|_{2}\|_{L^{q}}\|U_{11}\|_{L^{q}}$. Furthermore, Jensen’s inequality gives $\|\|Z_{1}\|_{2}\|_{L^{q}}=\|\|Z_{1}\|_{2}^{2}\|_{L^{q/2}}^{1/2}\geq\|\|Z_{1}\|_{2}^{2}\|_{L^{1}}^{1/2}=\sqrt{p}$. Combining the last several steps and the assumption that $\|\eta_{1}\|_{L^{q}}\lesssim\sqrt{p}$, we conclude that
$\|\langle v,X_{1}\rangle\|_{L^{q}}\lesssim\|w\|_{2}=\|\langle v,X_{1}\rangle\|_{L^{2}}^{2},$ (42)
@@
Regarding the density condition (ii), first note that $X_{1j}/\sigma_{j}$ is equal in distribution to $\eta_{1}Z_{11}/\|Z_{1}\|_{2}$ for every $j=1,\dots,p$. Also note that the assumption $\limsup_{t\to 0^{+}}\frac{1}{t}\mathbf{P}(\eta_{1}^{2}/p\leq t)\lesssim 1$ implies that $\eta_{1}$ does not have an atom at 0. Combining this with the fact that $Z_{11}/\|Z_{1}\|_{2}$ has a Lebesgue density [18, eqn. 1.26] and that $\eta_{1}$ is independent of $Z_{1}$, it is straightforward to check that the distribution of $\eta_{1}Z_{11}/\|Z_{1}\|_{2}$ is absolutely continuous with respect to Lebesgue measure, and thus has a Lebesgue density. To establish the boundedness of this density, the Lebesgue differentiation theorem ensures that it is sufficient to show
$\limsup_{t\to 0^{+}}\ \sup_{x\in\mathbb{R}}\ \frac{1}{t}\mathbf{P}\Big{(}\Big{|}\frac{\eta_{1}Z_{11}}{\|Z_{1}\|_{2}}-x\Big{|}\leq t\Big{)}\lesssim 1.$ (43)
Since the random variable $Z_{11}/\|Z_{1}\|_{2}$ has a symmetric unimodal distribution [18, eqn. 1.26], we may apply Anderson’s lemma [3] conditionally on $\eta_{1}$ to conclude that the probability in (43) is maximized when $x=0$. Furthermore, by splitting the probability along the events $\{\eta_{1}\leq\sqrt{pt}\}$ and $\{\eta_{1}>\sqrt{pt}\}$, we have
$\mathbf{P}\Big{(}\Big{|}\frac{\eta_{1}Z_{11}}{\|Z_{1}\|_{2}}-x\Big{|}\leq t\Big{)}\ \leq\ \mathbf{P}\Big{(}\frac{|Z_{11}|}{\|Z_{1}\|_{2}}\leq\frac{t}{\sqrt{pt}}\Big{)}+\mathbf{P}(\eta_{1}\leq\sqrt{pt}).$
By assumption, the second term on the right satisfies $\limsup_{t\to 0^{+}}\frac{1}{t}\mathbf{P}(\eta_{1}\leq\sqrt{pt})\lesssim 1$. Also, since there is an absolute constant $c>0$ such that $\mathbf{P}(Z_{11}^{2}\leq x)\leq cx$ holds for all $x>0$, it follows that for any $t\in(0,1)$,
$\begin{split}\frac{1}{t}\mathbf{P}\Big{(}\frac{|Z_{11}|}{\|Z_{1}\|_{2}}\leq\frac{t}{\sqrt{pt}}\Big{)}&\ =\ \frac{1}{t}\mathbf{P}\bigg{(}Z_{11}^{2}\leq\frac{t(Z_{12}^{2}+\cdots+Z_{1p}^{2})}{(1-t/p)p}\bigg{)}\\[5.69046pt] &\ \leq\ \frac{c}{t}\,\mathbf{E}\bigg{(}\frac{t(Z_{12}^{2}+\cdots+Z_{1p}^{2})}{(1-t/p)p}\bigg{)}\\[5.69046pt] &\ \leq\ \frac{c}{1-t},\end{split}$
which clearly implies (43).
Separable case. Suppose that $X_{1}$ is has a centered separable distribution so that $X_{1}=\Sigma^{1/2}\zeta_{1}$, where $\zeta_{1}=(\zeta_{11},\dots,\zeta_{1p})$ has i.i.d. entries with $\mathbf{E}(\zeta_{11})=0$, and $\textup{var}(\zeta_{11})=1$. To check the $L^{q}$-$L^{2}$ moment equivalence condition (i), it suffices to show that $\|\langle w,\zeta_{1}\rangle\|_{L^{q}}\lesssim\|\langle w,\zeta_{1}\rangle\|_{L^{2}}$ for any $w\in\mathbb{R}^{p}$. Using Rosenthal’s inequality (Lemma 16), and the assumption that $\max_{1\leq j\leq p}\|\zeta_{1j}\|_{L^{q}}\lesssim 1$, we have
@@
as needed. Finally, the density condition (ii) is a direct consequence of Theorem 1.2 in the paper [49] and the assumption that $\max_{1\leq j\leq p}\|g_{j}\|_{L^{\infty}}\lesssim 1$, where $g_{j}$ is the Lebesgue density of $\zeta_{1j}$. ∎
@@ Appendix H Background results
Lemma 16 (Rosenthal inequalities [25]).
@@
Lemma 17 (Lower-tail bound for Gaussian maxima [37]).
@@
The next result is a variant of Lemma A.7 in [52] that can be proven in essentially the same way.
@@
[…]
Lemma 19 (Bentkus’ Berry-Esseen Theorem [5]).
@@
Let $(\zeta_{1},\ldots,\zeta_{d})$ be a Gaussian random vector, and suppose the parameter $\underline{\sigma}^{2}:=\min_{1\leq j\leq d}\textup{var}(\zeta_{j})$ is positive. Then, for any fixed $\epsilon>0$,
@@
This version of Nazarov’s inequality appears in Lemma 4.3 of [10] and originates from [43].
@@
[…]
This statement of the Fuk-Nagaev inequality is based on [47, eqn. 1.7].
@@
[…]
The result above is the modification of Corollary A.6.3 in [55]. In that reference, the success probability $\mathbf{E}(\xi_{1})$ of the Bernoulli random variables is assumed to be less than $1/e$, but in the formulation above, the result holds for all values of $\mathbf{E}(\xi_{1})$.

### v2

@@ Appendix F Results on median-of-means estimators
Proof. The proof of Lemma 12 can be repeated with the i.i.d. random variables $\frac{1}{2}(X_{ij}-X_{i^{\prime}j})^{2}$, playing the role that $X_{ij}$ previously did. Also note that because $\textup{var}\big(\frac{1}{2}(X_{ij}-X_{i^{\prime}j})^{2}\big)\lesssim\sigma_{j}^{4}$ holds under Assumption 1((i)), the parameter $\sigma_{j}^{4}$ plays the role that $\sigma_{j}^{2}$ did in the context of Lemma 12. ∎
@@
Since the independent random variables $X_{1j}/\sigma_{j}$ and $X_{(1+\ell_{n}/2)j}/\sigma_{j}$ have densities whose $L^{\infty}$ norms are $\mathcal{O}(1)$ under Assumption 1((ii)), it follows from Young’s convolution inequality [stein1971introduction, p.178] that the random variable $\frac{1}{\sqrt{2}\sigma_{j}}(X_{1j}-X_{(\ell_{n}/2+1)j})$ also has a density whose $L^{\infty}$ norm is $\mathcal{O}(1)$, and so
@@
where the last step uses $\ell_{n}\asymp n/\log(n)$. Note also that the final bound $n^{-(2+\delta)}$ can be replaced with any fixed positive power of $n^{-1}$, but the current form is all that is needed. ∎
@@ Appendix G Proof of Proposition 1
Elliptical case. Suppose $X_{1}$ is a centered elliptical random vector of the form $X_{1}=\eta_{1}\Sigma^{1/2}Z_{1}/\|Z_{1}\|_{2}$, where $Z_{1}$ is a standard Gaussian $p$-dimensional Gaussian vector, and $\eta_{1}$ is independent of $Z_{1}$ with $\mathbf{E}(\eta_{1}^{2})=p$. We first check the $L^{q}$-$L^{2}$ moment equivalence condition (i) with $q=4+\delta$. Letting $w=\Sigma^{1/2}v$ for a generic vector $v\in\mathbb{R}^{p}$, a direct calculation gives
@@
Because the distribution of $U_{1}=Z_{1}/\|Z_{1}\|_{2}$ is invariant to orthogonal transformations, it follows that the random variables $\langle w,U_{1}\rangle$ and $\|w\|_{2}\langle e_{1},U_{1}\rangle$ are equal in distribution, where $e_{1}$ is the first standard basis vector. Therefore,
@@
The quantity $\|U_{11}\|_{L^{q}}$ is at most of order $1/\sqrt{p}$, which can be shown as follows. Due to the independence of $U_{11}$ and $\|Z_{1}\|_{2}$, we have $\|Z_{11}\|_{L^{q}}=\|\|Z_{1}\|_{2}\|_{L^{q}}\|U_{11}\|_{L^{q}}$. Furthermore, Lyapunov’s inequality gives $\|\|Z_{1}\|_{2}\|_{L^{q}}=\|\|Z_{1}\|_{2}^{2}\|_{L^{q/2}}^{1/2}\geq\|\|Z_{1}\|_{2}^{2}\|_{L^{1}}^{1/2}=\sqrt{p}$. Combining the last several steps and the assumption that $\|\eta_{1}\|_{L^{q}}\lesssim\sqrt{p}$, we conclude that
$\|\langle v,X_{1}\rangle\|_{L^{q}}\lesssim\|w\|_{2}=\|\langle v,X_{1}\rangle\|_{L^{2}},$ (41)
@@
Regarding the density condition (ii), note that if $e_{j}$ is the $j$th standard basis vector, then the vector $w=\Sigma^{1/2}e_{j}/\sigma_{j}$ satisfies $\|w\|_{2}=1$. So, the discussion above shows that all the random variables $X_{11}/\sigma_{1},\dots,X_{1p}/\sigma_{p}$ have the same distribution, which is that of $\eta_{1}\langle e_{1},U_{1}\rangle$. In particular, if the random variable $X_{11}/\sigma_{1}$ has a Lebesgue density $f_{1}$ such that $\|f_{1}\|_{L^{\infty}}\lesssim 1$, then condition (ii) holds, which completes the proof in the elliptical case.
Separable case. Suppose that $X_{1}$ has a centered separable distribution so that $X_{1}=\Sigma^{1/2}\zeta_{1}$, where $\zeta_{1}=(\zeta_{11},\dots,\zeta_{1p})$ has i.i.d. entries with $\mathbf{E}(\zeta_{11})=0$, and $\textup{var}(\zeta_{11})=1$. To check the $L^{q}$-$L^{2}$ moment equivalence condition (i), it suffices to show that $\|\langle w,\zeta_{1}\rangle\|_{L^{q}}\lesssim\|\langle w,\zeta_{1}\rangle\|_{L^{2}}$ for any $w\in\mathbb{R}^{p}$. Using Rosenthal’s inequality (Lemma 16), and the assumption that $\max_{1\leq j\leq p}\|\zeta_{1j}\|_{L^{q}}\lesssim 1$, we have
@@
as needed. Finally, the density condition (ii) is a direct consequence of Theorem 1.2 in the paper [rudelson2015small] and the assumption that $\max_{1\leq j\leq p}\|g_{j}\|_{L^{\infty}}\lesssim 1$, where $g_{j}$ is the Lebesgue density of $\zeta_{1j}$. ∎
@@ Appendix H Background results
Lemma 16 (Rosenthal inequalities [Rosenthal]).
@@
Lemma 17 (Lower-tail bound for Gaussian maxima [lopes2022sharp]).
@@
The next result is a variant of Lemma A.7 in [spokoiny2015bootstrap] that can be proven in essentially the same way.
@@
$\sup_{s\in\mathbb{R}}\bigg|\mathbf{P}\Big(\max_{1\leq j\leq d}\xi_{j}\leq s\Big)-\mathbf{P}\Big(\max_{1\leq j\leq d}\zeta_{j}\leq s\Big)\bigg|\ \leq\ 2\|A-I_{d}\|_{F}.$
Lemma 19 (Bentkus’ Berry-Esseen Theorem [Bentkus:2003]).
@@
Let $(\zeta_{1},\ldots,\zeta_{d})$ be a Gaussian random vector, and suppose the parameter $\underline{\sigma}^{2}\mathrel{\mathop{\ordinarycolon}}=\min_{1\leq j\leq d}\textup{var}(\zeta_{j})$ is positive. Then, for any fixed $\epsilon>0$,
@@
This version of Nazarov’s inequality appears in Lemma 4.3 of [chernozhukov2016empirical] and originates from [nazarov2003maximal].
@@
$\mathbf{P}\bigg(\Big|\sum_{i=1}^{n}\xi_{i}\Big|\geq s\bigg)\ \leq\ 2\Big(\frac{q+2}{qs}\Big)^{q}\sum_{i=1}^{n}\mathbf{E}|\xi_{i}|^{q}+2\exp\bigg(\frac{-2s^{2}}{(q+2)^{2}e^{q}\sum_{i=1}^{n}\mathbf{E}(\xi_{i}^{2})}\bigg).$
This statement of the Fuk-Nagaev inequality is based on [rio2017constants, eqn. 1.7].
@@
$\mathbf{P}\bigg(\frac{1}{n}\sum_{i=1}^{n}\xi_{i}\geq\frac{1}{2}\Big)\ \leq\ 2\big(e\mathbf{E}(\xi_{1})\big)^{n(1/2-\mathbf{E}(\xi_{1}))^{2}}.$
The result above is the modification of Corollary A.6.3 in [van1996weak]. In that reference, the success probability $\mathbf{E}(\xi_{1})$ of the Bernoulli random variables is assumed to be less than $1/e$, but in the formulation above, the result holds for all values of $\mathbf{E}(\xi_{1})$.